I need to factor completely, but forgot how.
8x^3-27^6

Answer by nerdybill(7384)


You're suppose to recognize that this is a "special factoring" case -- "difference of cubes":
a^3 + b^3 = (a + b)(a^2 – ab + b^2)
.
8x^3-27^6
can be rewritten as:
(2x)^3-(27^2)^3
Therefore,
a is 2x
b is 27^2
.
(2x)^3 + (27^2)^3 = (2x + 27^2)((2x)^2 – (2x)(27^2) + (27^2)^2)
You can proceed to simplify the above...or, you can leave it the way it is.